# Bibliography.bib

### v1

@@ 3.3.1. Nice subcomplexes
The meaning of this is that any $\Delta$ in the ring of $K^{\prime}$ can be thus seen as the join of two faces, $A$ and $B$, with $A\in K^{\prime}$ and $B$ disjoint from $K^{\prime}$. This is useful in order to interpolate from a map/section given over $A$ to a map/section given over $B$ (as in Section 4.2). The following provides a useful criterion for niceness.
@@
3.3.2. Nice covers via convexity
@@ 3.5. Generalized crystalline subdivision
For the following statement, to ease the notation, we will write $B(A,r)$ for the $r$-neighborhood of a subset $A\subset{\mathbb{R}}^{N}$.
@@
  • agrees with $K^{(i)}\cap B^{(i)}$,
  • is the barycentric cone off on the ring of $K^{(i)}\cap B^{(i)}$, and
  • agrees with $K_{\ell_{0}+i+1}$ outside of it
@@
We bound the maximum number of colors $C$ as follows. Each simplex simplex is involved in at most one coning off process. Hence we can take $C=C^{\prime}\cdot A$, where we defined $C^{\prime}$ at the beginning of this proof, and $A$ is the resulting number of top-dimensional subsimplices after applying the barycentric cone off to a simplex $\Delta$. We can bound $A$ by $m-1+2^{m-1}$, where $m$ is the dimension of $\Delta$. To see this, let $F$ be the face of $\Delta$ subdivided once in a crystalline manner. Then the subdivision of $F$ contains $2^{m-1}$ simplices, which are each coned off to the barycenter. All the other faces of $\Delta$, of which there are $m-1$, are also coned off to the barycenter. ∎
@@ 4.1. The linearization statement
Fix a simplicial complex $K$ of pure dimension. Let $s:|K|\to{\mathbb{R}}^{n}$ be a map that is piecewise smooth with respect to $K$. We introduce the notation $s^{\mathrm{lin}}_{\ell}$ for the linearization of $s$ with respect to the $\ell$th crystalline subdivision $K_{\ell}$. Then:
@@
Consider an $m$-simplex $\Delta\in(\Delta^{m})_{\ell}$. Let ${\mathcal{D}}\subset\operatorname{Hom}({\mathbb{R}}^{m},{\mathbb{R}}^{n})$ be the subspace of all derivatives $d_{y}s:{\mathbb{R}}^{m}\to{\mathbb{R}}^{n}$ with $y\in\Delta$. Then by Taylor’s approximation formula, it holds that the maximum distance $d_{\mathrm{max}}$ between two elements in ${\mathcal{D}}$ is $O(\operatorname{r_{\mathrm{max}}}(\Delta))$. This maximum exists because $\Delta$ and hence ${\mathcal{D}}$ is compact.
We also fix a vertex $v\in\Delta$ and $e_{1},\dots e_{m}\in{\mathbb{R}}^{m}$ such that the other vertices of $\Delta$ are $\{v+e_{i}\}_{i}$. Then $s_{\ell}^{\mathrm{lin}}|_{\Delta}$ is the map that sends $x\in\Delta$ to $s(v)+A(x-v)$ where $A$ is the matrix with columns $s(v+e_{i})-s(v)$. By the mean value theorem there exist elements $y_{ij}\in\Delta$ such that $A_{ij}=(d_{y_{ij}}s_{j})e_{i}$. Hence we see that for the $j$th row of $A$ we obtain $A_{j}x=(d_{v}s_{j})x+O(d_{\mathrm{max}}|x|)$ for any $x\in{\mathbb{R}}^{m}$.
@@ 4.2. Modifying maps on a face
Let $\Delta$ be a linear simplex given as the join of two opposing faces $A$ and $B$. Let $t:\Delta\to[0,1]$ be the join parameter: the affine function that is 00 over $A$ and $1$ over $B$. Suppose that $s_{A},s_{B}:\Delta\to{\mathbb{R}}^{n}$ are two smooth maps. We define the interpolation of $s_{A}$ and $s_{B}$ over $\Delta$ with respect to $A$ and $B$, denoted as $\operatorname{interpolate}_{\Delta}(A,B,s_{A},s_{B})$, to be the map $ts_{A}+(1-t)s_{B}$.
@@
Fix a linear $m$-simplex $\Delta$ spanned by two opposing faces $A,B$. Consider moreover three maps $u_{1},u_{2}:A\rightarrow{\mathbb{R}}^{n}$, $v:B\to{\mathbb{R}}^{n}$, and the corresponding interpolations $s_{i}=\operatorname{interpolate}_{\Delta}(A,B,u_{i},v)$. Then the following bounds hold:
@@ 4.2.1. The linear case
Fix a linear $m$-simplex $\Delta$ spanned by two opposing faces $A,B$. Consider moreover three linear maps $u_{1},u_{2}:A\rightarrow{\mathbb{R}}^{n}$, $v:B\to{\mathbb{R}}^{n}$, and the corresponding joins $s_{i}=\operatorname{join}_{\Delta}(A,B,v,u_{i})$. Then the following bounds hold:
@@
  • $d_{C^{1}}(s_{1},s_{2})=d_{C^{0}}(s_{1},s_{2})+O\left(\frac{d_{C^{0}}(u_{1},u_{2})}{\operatorname{r_{\mathrm{min}}}(\Delta)}\right)$.
@@
We define affine sections $\tilde{s}_{i}:\Delta^{m}\to{\mathbb{R}}^{n}$ by requiring that $\tilde{s}_{i}(0)=s_{i}(0)$ and $\tilde{s}_{i}(e_{i})=s_{i}(v_{i})$, where the vectors $e_{i}$ are the unit vectors in ${\mathbb{R}}^{m}$. If we let $T$ be the linear transformation $\Delta\to\Delta^{m}$ sending $v_{i}$ to $e_{i}$, we see that $s_{i}=\tilde{s}_{i}\circ T$. We also observe that the operator norm of $T$ is $\lVert T\rVert=\frac{1}{\operatorname{r_{\mathrm{min}}}(\Delta)}$.
@@
$d_{C^{0}}(ds_{1},ds_{2})$ $=\lVert D_{1}T-D_{2}T\rVert\leq\lVert D_{1}-D_{2}\rVert\lVert T\rVert=O\left(d_{C^{0}}(u_{1},u_{2})\cdot\dfrac{1}{\operatorname{r_{\mathrm{min}}}(\Delta)}\right).\qed$
@@ 4.3. Relative linearization
Given an open neighborhood ${\mathcal{O}p}(|K^{\prime}|)$ we can find a finite collection of convex opens $\{U_{i}\}_{i=1}^{I}$ in ${\mathbb{R}}^{N}$ so that $\cup_{i}U_{i}$ intersected with $|K|$ is a neighborhood of $|K^{\prime}|$ contained in ${\mathcal{O}p}(|K^{\prime}|)$. We fix such a collection and assume that $\ell$ is large enough so that each simplex in $K_{\ell}$ is either contained in one of the $U_{i}$ or is disjoint from $|K^{\prime}|$. By Lemma 3.9 it follows that each subcomplex $K^{i}\coloneqq U_{i}\cap K$ is nice.
We now construct maps $s_{i}$ by induction on $i$ which each satisfy the requirements of the statement on $\cup_{j=0}^{i}U_{i}\cap K^{\prime}$ instead of the whole of $K^{\prime}$. We set $s_{0}:=s$, so that the base case $i=1$ follows the same proof as the inductive step. Hence assume $s_{i-1}$ is constructed. Over $K^{i}$ we use Lemma 4.2 to linearize $s_{i-1}$, yielding some map $\overline{s}_{i}$. Over the ring of $K^{i}$ we apply Lemma 4.4 to interpolate between $\overline{s}_{i}$ and $s_{i-1}$; this is what we call $s_{i}$. We define $s^{\prime}_{\ell}$ as the end result of the induction, that is $s^{\prime}_{\ell}=s_{I}$.
@@ 5.2.1. Over a single simplex
Consider a linear simplex $\Delta$. The reasoning from the previous subsection states that, given $\varepsilon>0$ and a $1$-jet $\sigma_{1}$, we can choose $\sigma_{2}\in B_{f}(\sigma_{1},\varepsilon)$ and consider its linear extension $s_{1}$ over $\Delta$. Naturally, if we had chosen some other $\sigma_{2}\in B_{f}(\sigma,\varepsilon)$ instead of $\sigma_{1}$, we would have obtained a different linear extension $s_{2}$. However, the difference between $s_{1}$ and $s_{2}$ is bounded by $\varepsilon$ and the size of $\Delta$. The following is then immediate.
@@
We observe that the second claim would still hold if $\sigma_{1}$ and $\sigma_{2}$ had lain in the same fiber over $x\in M$, but the first one would not. To emphasize that $\sigma_{1}$ and $\sigma_{2}$ lie in the same front fiber, we refer to $s_{2}$ as an $\varepsilon$-slope-perturbation of $s_{1}$.
@@ 5.2.2. Over a color
We now extend this procedure to all the simplex of a given color. We package this into a definition:
@@
We depict this in Figure 5.
@@
Let $K$ be a colored simplicial complex of pure dimension $m$. Let $K^{\prime}$ be the collection of simplices of color $i$, assume it is finite. Suppose $s^{\prime}:|K|\to{\mathbb{R}}^{n}$ is the result of slope-perturbing a piecewise linear map $s:|K|\to{\mathbb{R}}^{n}$ by $\varepsilon$ over the $i$th color. Let $\operatorname{r_{\mathrm{max}}}$ be the maximum value of $\operatorname{r_{\mathrm{max}}}(\Delta)$ over $\operatorname{star}(K^{\prime})$ and let $\operatorname{r_{\mathrm{min}}}$ be the minimum value of $\operatorname{r_{\mathrm{min}}}(\Delta)$ over $\operatorname{star}(K^{\prime})$. Then, the following bounds hold:
@@
  • $d_{C^{1}}(s,s^{\prime})=O\left(\varepsilon\operatorname{r_{\mathrm{max}}}\left(1+\dfrac{\operatorname{r_{\mathrm{max}}}}{\operatorname{r_{\mathrm{min}}}}\right)\right)$.
@@ 6.1. Jiggling in the linear setting
Consider a finite simplicial complex $K$ of pure dimension, and an open and fiberwise dense differential relation :=${\mathcal{R}}:=\{{\mathcal{R}}_{\Delta}\}_{\Delta\in K^{(top)}}$. Then, given
@@
We write $C$ for the number of colors needed to color all the $K_{\ell}$ (Lemma 3.3). We write $K_{\ell}^{(\mathrm{top},i)}$ for the collection of top dimensional simplices in $K_{\ell}$ of color $i$.
@@
Choosing a good collection of constants. We now choose the constants $\{\varepsilon_{i}\}_{i\in[C]}$, $\{\delta_{i}\}_{i\in[C]}$ and $\ell$. A first requirement is that $\ell$ must be large enough so that $d_{C^{1}}(s,s_{\ell}^{\mathrm{lin}})<\varepsilon/2$, which can be achieved using Lemma 4.2.
According to Lemma 5.8 we can find $\varepsilon_{0}>0$ such that, for any subdivision $\ell$ and color $i$, the $C^{1}$-distance between $s$ and the result of slope-perturbing $s$ by $\varepsilon_{0}$ over the $i$th color in $K_{\ell}$ is at most $\frac{\varepsilon}{2C}$. In fact, we can assume that this $\varepsilon_{0}$ works for any piecewise smooth map $h:|K|\to{\mathbb{R}}^{n}$ that is $\varepsilon$-close to $s$ in the $C^{1}$-distance. To see this, two comments are relevant: First we note that this works for $s$ because according to Lemma 3.6, we can bound the size and the ratio of the size of any two simplices independently of $\ell$. Next, we point out that the $1$-jet of $h$ is contained in the $\varepsilon$-neighborhood of $j^{1}s$ in $J^{1}(|K|,{\mathbb{R}}^{n})$, which is compact, which allows us to find an $\varepsilon_{0}$ that works for all such $h$.
Corollary 5.5 then yields the constant $\delta_{0}>0$. Namely, it is chosen so that any $1$-jet $\sigma\in\overline{B(j^{1}s(|K|),\varepsilon)}$ admits an $\varepsilon_{0}$-slope-perturbation $\sigma^{\prime}$ such that the $\delta_{0}$-ball around it lies in ${\mathcal{R}}$. Moreover, the linear extension of each $\sigma^{\prime\prime}\in B(\sigma,\delta_{0})$ over a $\delta_{0}$-ball is a solution of ${\mathcal{R}}$.
We then keep going, arguing similarly, inductively on $i\in[C]$. Lemma 5.8 yields $\varepsilon_{i}>0$ small enough so that a $\varepsilon_{i}$-slope-perturbation over all simplices of the same color causes at most a $\delta_{i-1}/4$ perturbation over adjacent simplices, in a $C^{1}$-sense. Our choice of $\varepsilon_{i}$ then yields a $\delta_{i}>0$ using Corollary 5.5. We can assume without loss of generality that $\{\varepsilon_{i}\}_{i\in[C]}$ is a decreasing sequence and that $2\delta_{i+1}<\delta_{i}$ for all $i\in[C]$.
Lastly, we ask $\ell$ to be large enough so that $\max(1,||s||_{C^{1}}+2\varepsilon)\operatorname{r_{\mathrm{max}}}(\Delta)<\delta_{C}/2$ for every $\Delta\in K_{\ell}$. That this is possible follows from Lemma 3.6.
@@
By construction, item (IH2) is immediately satisfied. To see that item (IH1) is satisfied on each simplex $\Delta\in\cup_{j=0}^{i}K_{\ell}(i+1)$, we distinguish two cases: $\Delta$ is of color $i+1$ or of a previous color.
@@
where the latter term is bounded by $||s^{(i+1)}_{\ell}||_{C^{1}}\operatorname{r_{\mathrm{max}}}(\Delta)$. Our assumptions on $\ell$ thus imply that $\sigma^{\prime\prime}\in B(j^{1}s^{(i+1)}_{\ell}(v),\delta_{i+1})$. Hence, the section $\operatorname{LinExt}(\sigma^{\prime\prime})$ is a solution over $\Delta$.
Simplices of a previous color. Let now $\Delta$ be in $K_{\ell}^{(\mathrm{top},j)}$ for some $j<1+i$ and take take $\sigma^{\prime\prime}\in B(j^{1}s_{\ell}^{(i+1)}(x),\delta_{i+1}/2)$ with $x\in\Delta$. By definition of $\varepsilon_{i+1}$ it follows in particular that $d_{C^{0}}(j^{1}s_{\ell}^{(i)},j^{1}s_{\ell}^{(i+1)})<\delta_{i}/2$, and hence we see that
@@
  • $s^{\prime}$ is a piecewise solution of ${\mathcal{R}}$ on $|K|\setminus{\mathcal{O}p}(|K^{\prime\prime}|)$, and

### v2

@@ 3.3.1. Nice subcomplexes
We have illustrated Definition 3.9 in Figure 2. The meaning of niceness is that any $\Delta$ in the ring of $K^{\prime}$ can be thus seen as the join of two faces, $A$ and $B$, with $A\in K^{\prime}$ and $B$ disjoint from $K^{\prime}$. This is useful in order to interpolate from a map/section given over $A$ to a map/section given over $B$ (as in Section 4.2). The following provides a useful criterion for niceness.
@@
3.3.2. Obtaining nice subcomplexes
@@ 3.5. Generalized crystalline subdivision
For the following statement, to ease the notation, we will write $B(A,r)$ for the $r$-neighborhood of a subset $A\subset{\mathbb{R}}^{N}$. We also recall that we have defined when simplicial complexes agree in Definition 2.8.
@@
  • agrees with $K^{(i)}$ on $B^{(i)}$,
  • agrees with $K_{\ell_{0}+i+1}$ outside of $\operatorname{ring}(K^{(i)}\cap B^{(i)},K^{(i)})$, and
  • is the barycentric cone off on $\operatorname{ring}(K^{(i)}\cap B^{(i)},K^{(i)})$.
@@
We bound the maximum number of colors $C$ as follows. Each simplex is involved in at most one coning off process. Hence we can take $C=C^{\prime}\cdot A$, where we defined $C^{\prime}$ at the beginning of this proof, and $A$ is the resulting number of top-dimensional subsimplices after applying the barycentric cone off to a simplex $\Delta$. We can bound $A$ by $m-1+2^{m-1}$, where $m$ is the dimension of $\Delta$. To see this, let $F$ be the face of $\Delta$ subdivided once in a crystalline manner. Then the subdivision of $F$ contains $2^{m-1}$ simplices, which are each coned off to the barycenter. All the other faces of $\Delta$, of which there are m1m-1, are also coned off to the barycenter. ∎
@@ 4.1. The linearization statement
[…]
@@
Consider an $m$-simplex $\Delta\in(\Delta^{m})_{\ell}$. Let ${\mathcal{D}}\subset\operatorname{Hom}({\mathbb{R}}^{m},{\mathbb{R}}^{n})$ be the subspace of all derivatives $d_{y}s:{\mathbb{R}}^{m}\to{\mathbb{R}}^{n}$ with $y\in\Delta$. Then by Taylor’s approximation formula, it holds that the maximum distance $d_{\mathrm{max}}$ between two elements in $\mathcal{D}$} is $O(\operatorname{{r_{\mathrm{max}}}}(\Delta))$. This maximum exists because $\Delta$ and hence $\mathcal{D}$} is compact. We note that we obtain in particular for all $i\in\{1,\dots,m\}$ and $y_{1},y_{2}\in\Delta$ that
$|(d_{y_{1}}s_{j})e_{i}-(d_{y_{2}}s_{j})e_{i}|\leq|(d_{y_{1}}s_{j})-(d_{y_{2}}s_{j})|\leq d_{\mathrm{max}}.$
We also fix a vertex $v\in\Delta$ and $e_{1},\dots e_{m}\in{\mathbb{R}}^{m}$ such that the other vertices of $\Delta$ are $\{v+e_{i}\}_{i}$. Then $s_{\ell}^{\mathrm{lin}}|_{\Delta}$ is the map that sends $x\in\Delta$ to s(v)+A(xv)s(v)+A(x-v) where $A$ is the matrix with columns $s(v+e_{i})-s(v)$. By the mean value theorem there exist elements $y_{ij}\in\Delta$ such that $A_{ij}=(d_{y_{ij}}s_{j})e_{i}$. Hence we see that for the $j$th row of $A$ we obtain $A_{j}x=(d_{v}s_{j})x+O(d_{\mathrm{max}}|x|)$ for any $x\in{\mathbb{R}}^{m}$.
@@ 4.2. Modifying maps on a face
Let $\Delta$ be a linear simplex given as the join of two opposing faces $A$ and $B$. Let $t:\Delta\to[0,1]$ be the join parameter: the affine function that is 0 over $A$ and $1$ over $B$. Suppose that $s_{A},s_{B}:\Delta\to{\mathbb{R}}^{n}$ are two smooth maps. We define the interpolation of $s_{A}$ and $s_{B}$ over $\Delta$ with respect to $A$ and $B$, denoted as $\operatorname{interpolate}_{\Delta}(A,B,s_{A},s_{B})$, to be the map $x\mapsto t(x)s_{A}(x)+(1-t(x))s_{B}(x)$.
@@
Fix a linear $m$-simplex $\Delta$ spanned by two opposing faces $A,B$. Consider moreover three maps $u_{1},u_{2}:A\rightarrow{\mathbb{R}}^{n}$, $v:B\to{\mathbb{R}}^{n}$, and the interpolations $s_{i}=\operatorname{interpolate}_{\Delta}(A,B,u_{i},v)$. Then the following bounds hold:
@@ 4.2.1. The linear case
Fix a linear $m$-simplex $\Delta$ spanned by two opposing faces $A,B$. Consider moreover the linear maps $v:A\to{\mathbb{R}}^{n}$ and $u_{1},u_{2}:B\rightarrow{\mathbb{R}}^{n}$, and the corresponding joins $s_{i}=\operatorname{join}_{\Delta}(A,B,v,u_{i})$. Then the following bounds hold:
@@
  • $d_{C^{1}}(s_{1},s_{2})=d_{C^{0}}(s_{1},s_{2})+O\left(d_{C^{0}}(u_{1},u_{2})\cdot{\Lambda}(\Delta)\right)$.
@@
We define affine sections $\tilde{s}_{i}:\Delta^{m}\to{\mathbb{R}}^{n}$ by requiring that $\tilde{s}_{i}(0)=s_{i}(0)$ and $\tilde{s}_{i}(e_{i})=s_{i}(v_{i})$, where the vectors $e_{i}$ are the unit vectors in ${\mathbb{R}}^{m}$. If we let $T$ be the linear transformation $\Delta\to\Delta^{m}$ sending $v_{i}$ to $e_{i}$, we see that $s_{i}=\tilde{s}_{i}\circ T$. We recall from Lemma 3.7 that the operator norm of $T$ is $O({\Lambda}(\Delta))$.
@@
$d_{C^{0}}(ds_{1},ds_{2})$ $=O\left(\lVert D_{1}T-D_{2}T\rVert\right)$
$\leq O\left(\lVert D_{1}-D_{2}\rVert\lVert T\rVert\right)=O\left(d_{C^{0}}(u_{1},u_{2})\cdot{\Lambda}(\Delta)\right).\qed$
@@ 4.3. Relative linearization
Given an open neighborhood ${\mathcal{O}p}(|K^{\prime}|)$ we can find a finite collection of convex and open subsets $\{U_{i}\}_{i=1}^{I}$ of ${\mathbb{R}}^{N}$ so that $\cup_{i}U_{i}$ intersected with $|K|$ is a neighborhood of $|K^{\prime}|$ contained in ${\mathcal{O}p}(|K^{\prime}|)$. We fix such a collection and assume that $\ell$ is large enough so that each simplex in $K_{\ell}$ is either contained in one of the $U_{i}$ or is disjoint from $|K^{\prime}|$. By Lemma 3.11 it follows that each subcomplex $K^{i}\coloneqq U_{i}\cap K$ is nice.
We now construct maps $s_{i}:|K|\to{\mathbb{R}}^{n}$ by induction on $i$ which each satisfy all of the requirements of the statement, as long as we restrict $K^{\prime}$ to $\cup_{j=0}^{i}U_{j}\cap K^{\prime}$. We set $s_{0}:=s$, so that the base case $i=1$ follows the same proof as the inductive step. Hence assume $s_{i-1}$ is constructed. We use Lemma 4.2 to linearize $s_{i-1}$ over $|K^{i}|$, yielding some map $\overline{s}_{i}:|K^{i}|\to{\mathbb{R}}^{n}$. Over the ring of $K^{i}$ we interpolate between $\overline{s}_{i}$ and $s_{i-1}$; we call this interpolation $s_{i}$ and bound its difference to $s_{i-1}$ over the ring of $K^{i}$ using Lemma 4.4. We define $s^{\prime}_{\ell}$ as the end result of the induction, that is $s^{\prime}_{\ell}=s_{I}$.
@@ 5.2.1. Over a single simplex
Consider a linear simplex $\Delta$. The reasoning from the previous subsection states that, given $\varepsilon>0$ and a $1$-jet $\sigma_{1}\in J^{1}(\Delta,{\mathbb{R}}^{n})$, we can choose $\sigma_{2}\in B_{f}(\sigma_{1},\varepsilon)$ and consider its linear extension $s_{1}$ over $\Delta$. Naturally, if we had chosen some other $\sigma_{2}\in B_{f}(\sigma,\varepsilon)$ instead of $\sigma_{1}$, we would have obtained a different linear extension $s_{2}$. However, the difference between $s_{1}$ and $s_{2}$ is bounded by $\varepsilon$ and the size of $\Delta$. The following is then immediate.
@@
We observe that the second claim would still hold if $\sigma_{1}$ and $\sigma_{2}$ had lain in the same fiber over $x\in M$. The first claim would not, as we then have that $d_{C^{0}}(\sigma_{1},\sigma_{2})=O(\varepsilon+\varepsilon\operatorname{{r_{\mathrm{max}}}}(\Delta))$. To emphasize that $\sigma_{1}$ and $\sigma_{2}$ lie in the same front fiber, we refer to $s_{2}$ as an $\varepsilon$-slope-perturbation of $s_{1}$.
@@ 5.2.2. Over a color
We now extend this procedure to all the simplices of a given color. We package this into a definition:
@@
We note that $s^{\prime}$ is in particular piecewise linear and have depicted $s^{\prime}$ in Figure 5.
@@
Let $K$ be a colored simplicial complex of pure dimension $m$. Let $K^{\prime}$ be the collection of simplices of color $i$, assume it is finite. Suppose $s^{\prime}:|K|\to{\mathbb{R}}^{n}$ is the result of slope-perturbing a piecewise linear map $s:|K|\to{\mathbb{R}}^{n}$ by $\varepsilon$ over the $i$th color. Let $\operatorname{{r_{\mathrm{max}}}}$ and $\Lambda$} be the maximum value of respectively $\operatorname{{r_{\mathrm{max}}}}(\Delta)$ and ${\Lambda}(\Delta)$ over $\operatorname{star}(K^{\prime})$ and let $\operatorname{{r_{\mathrm{min}}}}$ be the minimum value of $\operatorname{{r_{\mathrm{min}}}}(\Delta)$ over $\operatorname{star}(K^{\prime})$. Then, the following bounds hold:
@@
  • $d_{C^{1}}(s,s^{\prime})=O\left(\varepsilon\operatorname{{r_{\mathrm{max}}}}\left(1+{\Lambda}\right)\right)$.
@@ 6.1. Jiggling in the linear setting
Consider a finite simplicial complex $K$ of pure dimension, and an open and fiberwise dense differential relation ${\mathcal{R}}:=\{{\mathcal{R}}_{\Delta}\}_{\Delta\in K^{(top)}}$, where ${\mathcal{R}}_{\Delta}\subset J^{1}(\Delta,{\mathbb{R}}^{n})$ for all $\Delta\in K^{(top)}$. Then, given
@@
We write $C$ for the number of colors needed to color all the $K_{\ell}$ (Lemma 3.3). We write $K_{\ell}^{(\mathrm{top},i)}$ for the collection of top-dimensional simplices in $K_{\ell}$ of color $i$.
@@
Choosing a good collection of constants. We now choose the sequences $\{\varepsilon_{i}\}_{i\in[C]}$ and $\{\delta_{i}\}_{i\in[C]}$ and $\ell$. A first requirement is that $\ell$ must be large enough such that $d_{C^{1}}(s,s_{\ell}^{\mathrm{lin}})<\varepsilon/2$, which can be achieved using Lemma 4.2.
We then claim that there exist sequences $\{\varepsilon_{i}\}_{i\in[C]}$ and $\{\delta_{i}\}_{i\in[C]}$ such that
  1. (1)
    an $\varepsilon_{i}$-slope-perturbation, of any piecewise linear map $h:|K|\to{\mathbb{R}}^{n}$ that is $\varepsilon$-close to $s$ in the $C^{1}$-distance, over all simplices of the same color, causes at most a $\delta_{i-1}/4$ perturbation over adjacent simplices, in a $C^{1}$-sense,
  2. (2)
    any $1$-jet $\sigma\in\overline{B(j^{1}s(|K|),\varepsilon_{i})}$ admits an $\varepsilon_{i}$-slope-perturbation $\sigma^{\prime}$ satisfying that the $\delta_{i}$-ball around it lies in $\mathcal{R}$}. Moreover, the linear extension of each $\sigma^{\prime\prime}\in B(\sigma^{\prime},\delta_{i})$ over the $\delta_{i}$-ball around $\pi(\sigma^{\prime\prime})$ is a solution of $\mathcal{R}$},
  3. (3)
    $\varepsilon_{i+1}<\varepsilon_{i}$, and
  4. (4)
    $2\delta_{i+1}<\delta_{i}$.
These sequences can be constructed inductively on $i\in[C]$. When we set $\delta_{-1}=\frac{\varepsilon}{2C}$ and $\varepsilon_{-1}=1$, we can treat the base case $i=0$ as we treat the inductive step. Hence we assume that we have constructed $\delta_{i-1}$ and $\varepsilon_{i-1}$. For the existence of $\varepsilon_{i}$, two comments are relevant: First we note that we can find $\varepsilon_{i}$ such that item (1) holds for $s^{\mathrm{lin}}_{\ell}$ because according to Lemma 3.8, we can bound the size of any simplex in $K_{\ell}$ in such a way that we can bound the $C^{1}$-distance independently of $\ell$. Next, we point out that the $1$-jet of $h$ is contained in the $\varepsilon$-neighborhood of $j^{1}s$ in $J^{1}(|K|,{\mathbb{R}}^{n})$, which is compact, which allows us to find an $\varepsilon_{0}$ that works for all such $h$. The constants $\delta_{i}$ such that item (2) holds, are produced by Corollary 5.5.
[…]
@@
By our definition of $\{\varepsilon_{i}\}_{i\in[C]}$, item (IH2) is immediately satisfied for $s_{\ell}^{(i)}$ and $s_{\ell}^{(+1i)}$. To see that item (IH1) is satisfied on each simplex $\Delta\in\cup_{j=0}^{i}K_{\ell}^{(i+1)}$, we distinguish two cases: $\Delta$ is of color $i+1$ or of a previous color.
@@
where the latter term is bounded by $s^{(i+1)}_{\ell}||_{C^{1}}\operatorname{{r_{\mathrm{max}}}}(\Delta)$. Our assumptions on $\ell$ thus imply that $\sigma^{\prime\prime}\in B(j^{1}s^{(i+1)}_{\ell}(v),\delta_{i+1})$. Hence, the section $\operatorname{LinExt}(\sigma^{\prime\prime})$ is a solution over $\Delta$.
Simplices of a previous color. Let now $\Delta$ be in $K_{\ell}^{(\mathrm{top},j)}$ for some $j<1+i$ and take take $\sigma^{\prime\prime}\in B(j^{1}s_{\ell}^{(i+1)}(x),\delta_{i+1}/2)$ with $x\in\Delta$. By definition of $\varepsilon_{i+1}$ it follows in particular that we have $d_{C^{0}}(j^{1}s_{\ell}^{(i)},j^{1}s_{\ell}^{(i+1)})<\delta_{i}/2$, and hence we see that
@@
  • $s^{\prime}$ is a piecewise smooth solution of $\mathcal{R}$} on $|K|\setminus{\mathcal{O}p}(|K^{\prime\prime}|)$, and